Given the circuit of capacitors and battery, determine the charge that forms on the 30uFcapacitor


30 uF 10 uF 20 V 2 ur I 40 uF
Answer #1

let
C1 = 30 micro F
C2 = 2 micro F
C3 = 10 micro F
C4 = 40 micro F

C3 and C4 are in series.
so, C34 = C3*C4/(C3 + C4)

= 10*40/(10 + 40)

= 8 micro F

the above combination is in parallel with C2.

so, C234 = C2 + C34

= 2 + 8

= 10 micro F

the above combination is in series with C1.

so, equivalent capacitance, Ceq = C1*C234/(C1 + C234)

= 30*10/(30 + 10)

= 7.50 micro F

charge on C1(30 micro F) capacitor, C1 = Ceq*V_battery

= 7.5*20

= 150 micro C (or) 1.50*10^-4 C <<<<<<<<------------Answer